OpenStudy (unklerhaukus):

a luminosity of 5 inverse femtobarns?! @Mathematics

OpenStudy (unklerhaukus):

how many inverse sheds/outhouses is that

OpenStudy (perl):

that sounds made up . are you doing a joke

OpenStudy (unklerhaukus):

CMS experiment at CERN has reported this data, it is a measure of collisions

OpenStudy (perl):

an outhouse is where you take a dump , is it not?

OpenStudy (unklerhaukus):

ha ha , http://en.wikipedia.org/wiki/Barn_%28unit%29

OpenStudy (perl):

so whats the whole question in its entirety. do not omit anything or you will be punished severely

OpenStudy (perl):

whats the question , im missing something here

OpenStudy (perl):

whats an inverse femtobarn, the inverse part ?

OpenStudy (perl):

The "inverse femtobarn" (fb−1) is a measurement of particle collision events per femtobarn. One inverse femtobarn is equal to around 70 million million (70 x 1012) collisions

OpenStudy (perl):

70*10^12 collisions , so if you want 5 femtobarns, you have 5*70*10^12 particle collision events per femtobarn

OpenStudy (perl):

A barn is the unit used by nuclear physicist for the size of an atomic nucleus of uranium. The cross-sectional area of a uranium nucleus is about 10 -24 square centimeters, small on the human scale, but large compared with other atomic nuclei. "Femto" means a factor of 10-15: a thousandth of a millionth of a millionth. A femtobarn, then, is 10-39 square centimeters ­ an incomprehensibly small unit of area.

OpenStudy (unklerhaukus):

ill pose the question: How many inverse outhouses are equivalent to 5 inverse femtobarns

OpenStudy (perl):

ok whats an outhouse

OpenStudy (perl):

an outhouse is 10^-34 m^2

OpenStudy (perl):

a femtobarn is 10^-43

OpenStudy (unklerhaukus):

right, so 1 outhouse 1x10^9 = 1 femtobarn?

OpenStudy (perl):

10^9 femtobarns = 1 outhouse

OpenStudy (unklerhaukus):

you got it the wrong way

OpenStudy (perl):

no i didnt , an outhouse is bigger than a femtobarn

OpenStudy (perl):

so 10^9 * 1 femtobarn = 1 outhouse

OpenStudy (perl):

a femtobarn is 10^-43 which is nine orders of magnitude smaller than an outhouse

OpenStudy (unklerhaukus):

so an outhouse is larger than a femtobarn

OpenStudy (perl):

yes it is , by nine orders of magntidue

OpenStudy (perl):

magnitude*

OpenStudy (perl):

an outhouse is a 'micro' barn, if that helps. and femto barn , is femto

OpenStudy (perl):

micro means 10^-6 of something, like a micro meter is 10^-6 meter. so if outhouse = microbarn, then it is 10^-6 of a barn

OpenStudy (perl):

outhouse = microbarn = 10^-6 *10^-28 = 10^-34

OpenStudy (unklerhaukus):

good

OpenStudy (perl):

interesting pic

OpenStudy (unklerhaukus):

ay? <====that pic?

OpenStudy (perl):

yep

OpenStudy (unklerhaukus):

well i find it interesting that you find it interesting

OpenStudy (perl):

yes , are you stoned there

OpenStudy (perl):

hehe

OpenStudy (perl):

what class is this femtobarn business

OpenStudy (unklerhaukus):

this question is not from a class, i am trying to comprehend the latest result from the CMS experiment at CERN (European Organization for Nuclear Research)